# Heavy-Tailed Linear Bandits: Huber Regression with One-Pass Update

### v1

@@ Abstract
We study the stochastic linear bandits with heavy-tailed noise. Two principled strategies for handling heavy-tailed noise, truncation and median-of-means, have been introduced to heavy-tailed bandits. Nonetheless, these methods rely on specific noise assumptions or bandit structures, limiting their applicability to general settings. The recent work (Huang et al., 2024) develops a soft truncation method via the adaptive Huber regression to address these limitations. However, their method suffers undesired computational cost: it requires storing all historical data and performing a full pass over these data at each round. In this paper, we propose a one-pass algorithm based on the online mirror descent framework. Our method updates using only current data at each round, reducing the per-round computational cost from $\mathcal{O}(t\log T)$ to $\mathcal{O}(1)$ with respect to current round $t$ and the time horizon $T$, and achieves a near-optimal and variance-aware regret of order $\widetilde{\mathcal{O}}\big{(}dT^{\frac{1-\varepsilon}{2(1+\varepsilon)}}\sqrt{\sum_{t=1}^{T}\nu_{t}^{2}}+dT^{\frac{1-\varepsilon}{2(1+\varepsilon)}}\big{)}$ where $d$ is the dimension and $\nu_{t}^{1+\varepsilon}$ is the $(1+\varepsilon)$-th central moment of reward at round $t$.
@@ 1 Introduction
Stochastic Linear Bandits (SLB) models sequential decision-making process with linear structured reward distributions, which has been extensively studied in the literature (Dani et al., 2008; Abbasi-Yadkori et al., 2011; Li et al., 2021). Specifically, in general SLB the observed reward at time $t$ is the inner product of the arm’s feature vector $X_{t}$ and an unknown parameter $\theta_{*}$, corrupted by sub-Gaussian noise $\eta_{t}$, namely,
@@
However, in many real-world scenarios, the noise in data often follows a heavy-tailed distribution, such as in financial markets (Cont and Bouchaud, 2000; Hull, 2012) and online advertising (Choi et al., 2020; Jebarajakirthy et al., 2021). This motivates the study of heavy-tailed linear bandits (HvtLB) model (Medina and Yang, 2016; Shao et al., 2018; Xue et al., 2020; Huang et al., 2024), where the noise $\eta_{t}$ in (1) satisfies that for some $\varepsilon\in(0,1],~{}\nu>0$,
@@
To address heavy-tailed noise, statistical methods in estimation and regression often rely on two key principles: truncation and median-of-means (MOM) (Lugosi and Mendelson, 2019). Truncation methods handle outliers by directly removing extreme data points. In the case of soft version truncation, such as Catoni’s M-estimator (Catoni, 2012), which reduces the impact of outliers by assigning them lower weights. This ensures robust estimates while maintaining potentially valuable information from extreme data points. Median-of-means (MOM) methods take a different approach by dividing the dataset into several groups, calculating the mean within each group, and then using the median of these group means as the final estimate. This method limits the influence of outliers to only a few groups, preventing them from affecting the overall dataset.
Table 1: Comparisons of our regret bounds and computational complexity to previous best-known results for heavy-tailed linear bandits. For the regret, the logarithmic dependence over $T$ is hidden by $\widetilde{\mathcal{O}}(\cdot)$-notation; $d$ is the dimension, $\nu_{t}$ is the moment bound. For the computational cost, we only keep the dependence on the time step $t$ and overall time step $T$.
@@
Huber HEAVY-OFUL (Huang et al., 2024) $\widetilde{\mathcal{O}}\left(dT^{\frac{1-\varepsilon}{2(1+\varepsilon)}}\sqrt{\sum_{t=1}^{T}\nu_{t}^{2}}+dT^{\frac{1-\varepsilon}{2(1+\varepsilon)}}\right)$ $\mathcal{O}(t\log T)$ instance-dependent bound
@@
For HvtLB, Shao et al. (2018) constructed algorithms using truncation and MOM-based least squares approach for parameter estimation, and achieved an $\widetilde{\mathcal{O}}\big{(}dT^{\frac{1}{1+\varepsilon}}\big{)}$ regret bound which is proven to be optimal. However, these approaches have some notable issues: truncation methods require the assumption of bounded absolute moments for rewards, which cannot vanish in deterministic case, making them suboptimal in noiseless scenarios; MOM methods rely on repeated arm pulling and assumption of a fixed arm set, which is only feasible in bandit settings. These limitations reveal that methods based on truncation and MOM heavily depend on specific assumptions or the bandit structure, making them challenging to adapt to broader scenarios, such as online MDPs and adaptive control. A more detailed discussion of these challenges is provided in Section 5.
Recently, Sun et al. (2020) made remarkable progress in the study of heavy-tailed statistics by proposing adaptive Huber regression, which leverages the Huber loss (Huber, 1964) to achieve non-asymptotic guarantees. Li and Sun (2023) further applied this technique to SLB, achieving an optimal and variance-aware regret bound under bounded variance conditions ($\varepsilon=1$). Subsequently, Huang et al. (2024) extended the adaptive Huber regression to handle heavy-tailed scenarios ($\varepsilon\in(0,1]$), achieving an optimal and instance-dependent regret bound of $\widetilde{\mathcal{O}}\big{(}dT^{\frac{1-\varepsilon}{2(1+\varepsilon)}}\sqrt{\sum_{t=1}^{T}\nu_{t}^{2}}+dT^{\frac{1-\varepsilon}{2(1+\varepsilon)}}\big{)}$ for HvtLB. Notably, the adaptive Huber regression does not rely on additional noise assumptions or the repeated pulling of bandit settings. This generality allows the approach to be further adapted to other decision-making scenarios, such as reinforcement learning with function approximation.
@@
where $\Theta$ is the feasible set of the parameter, $\lambda$ is the regularization parameter, and $\ell_{s}(\theta)$ represents the Huber loss at round $s$. Unlike the OFUL algorithm (Abbasi-Yadkori et al., 2011) in SLB with sub-Gaussian noise, which uses least squares (LS) for estimation and can update recursively based on the closed-form solution, the Huber loss is partially quadratic and partially linear. Solving (3) requires storing all historical data and performing a full pass over all historical data at each round $t$ to update the parameter estimation. This results in a per-round storage cost of $\mathcal{O}(t)$ and computational cost of $\mathcal{O}(t\log T)$ with respect to the current round $t$ and the time horizon $T$, making it computationally infeasible for large-scale or real-time applications. Thus, we need a one-pass algorithm for HvtLB that is as efficient as OFUL in SLB with sub-Gaussian noise, updating the estimation at each round $t$ using only the current data, without storing historical data. While one-pass algorithms based on the truncation and MOM approach have been proposed by Xue et al. (2023), but as noted earlier, these methods heavily rely on specific assumptions or the bandit structure. Given the attractive properties of Huber loss-based method (Huang et al., 2024), a natural question arises: Is it possible to design a one-pass Huber-loss-based algorithm that achieves optimal regret?
@@
Our Results.   In this work, we propose a Huber loss-based one-pass algorithm Hvt-UCB utilizing the Online Mirror Descent (OMD) framework, which achieves the optimal regret bound of $\widetilde{\mathcal{O}}(dT^{\frac{1}{1+\varepsilon}})$ while eliminating the need to store historical data. Additionally, our algorithm can further achieve the instance-dependent regret bound of $\widetilde{\mathcal{O}}\big{(}dT^{\frac{1-\varepsilon}{2(1+\varepsilon)}}\sqrt{\sum_{t=1}^{T}\nu_{t}^{2}}+dT^{\frac{1-\varepsilon}{2(1+\varepsilon)}}\big{)}$, where $\nu_{t}^{1+\varepsilon}$ is the $(1+\varepsilon)$-th central moment of reward at round $t$. Our approach preserves both optimal and instance-dependent regret guarantees achieved by Huang et al. (2024) while significantly reducing the per-round computational cost from $\mathcal{O}(t\log T)$ to $\mathcal{O}(1)$ with respect to the current round $t$ and the time horizon $T$. Moreover, our algorithm does not rely on any additional assumptions, making it more broadly applicable. A comprehensive comparison of our results and previous works on heavy-tailed linear bandits is presented in Table 1.
@@
Techniques.   Our main technical contribution is adapting the OMD framework, originally developed for regret minimization in online optimization, to parameter estimation in HvtLB. This adaptation requires linking the parameter gap (estimation error) with the loss value gap (regret) and involves two key components: (i) handling heavy-tailed noiseZhang et al. (2016) have adopted a variant of Online Newton Step (ONS) to Logistic Bandits settings, but their method cannot be applied to HvtLB since their method relies on the global strong convexity of the logistic loss, whereas the Huber loss is partially linear. We generalize their approach to the OMD framework and introduce a recursive normalization factor that ensures normal data stays in the quadratic region of the Huber loss with high probability, maintaining robustness in estimation. This normalization factor also introduces a negative term in the estimation error, which is crucial for controlling stability; (ii) supporting one-pass update: The OMD framework introduces a stability term in estimation error, which represents the impact of online updates. The adaptive normalization factor adds a multiplicative bias to this term, resulting in a positive $\widetilde{\mathcal{O}}(\sqrt{T})$ term that weakens the bound. This requires a careful design of the normalization factor and learning rate of OMD to cancel out the positive term using the negative term, ensuring stability and maintaining a tight bound.
Notations.   For a vector $\mathbf{x}\in\mathbb{R}^{d}$ and postive semi-definite matrix $V\in\mathbb{R}^{d\times d}$, we denote $\left\|\mathbf{x}\right\|_{V}=\sqrt{\mathbf{x}^{\top}V\mathbf{x}}$. For a strictly convex and continuously differentiable function $\psi:\mathcal{W}\rightarrow\mathbb{R}$, the Bregman divergence is define as $\mathcal{D}_{\psi}(\bm{w}_{1},\bm{w}_{2})=\psi(\bm{w}_{1})-\psi(\bm{w}_{2})-\left\langle\nabla\psi(\bm{w}_{2}),\bm{w}_{1}-\bm{w}_{2}\right\rangle$.
@@ 2 Related Work
Heavy-Tailed Bandits.   The multi-armed bandits (MAB) problem with heavy-tailed rewards, characterized by finite $(1+\varepsilon)$-moment, was first studied by Bubeck et al. (2013). They introduced two widely used robust estimation and regression methods from statistics into the heavy-tailed bandits setting: truncation and median-of-means (MOM). Medina and Yang (2016) extended these methods to heavy-tailed linear bandits (HvtLB) and proposed two algorithms: the truncation-based CRT algorithm and the MOM-based MoM algorithm, achieving $\widetilde{\mathcal{O}}\big{(}dT^{\frac{2+\varepsilon}{2(1+\varepsilon)}}\big{)}$ and $\widetilde{\mathcal{O}}\big{(}dT^{\frac{1+2\varepsilon}{1+3\varepsilon}}\big{)}$ regret bounds, respectively. Later, Shao et al. (2018) established an $\Omega\big{(}dT^{\frac{1}{1+\varepsilon}}\big{)}$ lower bound for HvtLB and introduced the truncation-based TOFU algorithm and the MOM-based MENU algorithm, both achieving an $\widetilde{\mathcal{O}}\big{(}dT^{\frac{1}{1+\varepsilon}}\big{)}$ regret bound. For HvtLB with finite arm set, Xue et al. (2020) established an $\Omega\big{(}d^{\frac{\varepsilon}{1+\varepsilon}}T^{\frac{1}{1+\varepsilon}}\big{)}$ lower bound, and proposed the truncation-based BTC algorithm and MOM-based BMM algorithm, both of which achieve an $\widetilde{\mathcal{O}}\big{(}\sqrt{d}T^{\frac{1}{1+\varepsilon}}\big{)}$ regret bound, reducing the dependence on the dimension $d$ under the finite arm set setting. To the best of our knowledge, the two works (Kang and Kim, 2023; Li and Sun, 2023) first introduce Huber loss to linear bandits. Specifically, Kang and Kim (2023) studied the heavy-tailed linear contextual bandits (HvtLCB) with fixed and finite arm set, where the context $X_{t}$ at each round is independently and identically distributed from a fixed distribution. They proposed the Huber-Bandit algorithm based on Huber regression, achieving an $\widetilde{\mathcal{O}}\big{(}\sqrt{d}T^{\frac{1}{1+\varepsilon}}\big{)}$ regret bound. On the other hand, Li and Sun (2023) were the first to apply adaptive Huber regression technique (Sun et al., 2020) to SLB, focusing on noise with finite variance (namely, HvtLB with $\varepsilon=1$). They proposed AdaOFUL algorithm, achieving an optimal variance-aware regret of $\widetilde{\mathcal{O}}(d\sqrt{\sum_{t=1}^{T}\nu_{t}^{2}})$, where $\nu_{t}$ is the upper bound of the variance. This result is subsequently improved by Huang et al. (2024), which is also the most related work to ours. Huang et al. (2024) considered general heavy-tailed scenarios with $\varepsilon\in(0,1]$ and proposed HEAVY-OFUL with an instance-dependent regret of $\widetilde{\mathcal{O}}\big{(}dT^{\frac{1-\varepsilon}{2(1+\varepsilon)}}\sqrt{\sum_{t=1}^{T}\nu_{t}^{2}}+dT^{\frac{1-\varepsilon}{2(1+\varepsilon)}}\big{)}$ where $\nu_{t}^{1+\varepsilon}$ is the $(1+\varepsilon)$-th central moment of reward at round $t$. We emphasize that all the above Huber regression-based algorithms for heavy-tailed bandits are not updated in a one-pass manner.
@@
One-Pass Bandit Learning.   In SLB, Abbasi-Yadkori et al. (2011) proposed the OFUL algorithm, achieving an $\mathcal{O}(d\sqrt{T})$ regret bound. It uses Least Squares (LS) for parameter estimation, naturally supporting one-pass updates with a closed-form solution. As an extension of SLB, Generalized Linear Bandits (GLB) were first introduced by Filippi et al. (2010). Their proposed GLM-UCB algorithm achieves an $\mathcal{O}(\kappa^{-1}d\sqrt{T})$ regret bound, where $\kappa$ measures the non-linearity of the generalized linear model. This algorithm relies on Maximum Likelihood Estimation (MLE) which doesn’t have a closed-form solution like LS in SLB and requires storing all historical data, resulting in a per-round computational cost of $\mathcal{O}(t)$. To this end, Zhang et al. (2016) proposed OL2M, the first one-pass algorithm for Logistic Bandits (a specific class of GLB), achieving an $\mathcal{O}(\kappa^{-1}d\sqrt{T})$ regret bound with per-round computational cost of $\mathcal{O}(1)$. Later, Jun et al. (2017) explored one-pass algorithms for GLB by introducing the online-to-confidence-set conversion. Using this conversion, they developed GLOC, a one-pass algorithm that achieves the same theoretical guarantees and per-round cost as OL2M but is applicable to all GLB. Xue et al. (2023) considered heavy-tailed GLB and proposed two one-pass algorithms: truncation-based CRTM and mean-of-medians-based CRMM, both achieving an $\widetilde{\mathcal{O}}\big{(}\kappa^{-1}dT^{\frac{1}{1+\varepsilon}}\big{)}$ regret bound. Subsequent works have tackled more challenging topics, such as reducing the dependence on $\kappa$ with one-pass algorithms or designing one-pass algorithms for more complex models like Multinomial Logistic Bandits and Multinomial Logit MDPs (Zhang and Sugiyama, 2023; Li et al., 2024; Lee and Oh, 2024; Li et al., 2025).

### v2

@@ Abstract
We study the stochastic linear bandits with heavy-tailed noise. Two principled strategies for handling heavy-tailed noise, truncation and median-of-means, have been introduced to heavy-tailed bandits. Nonetheless, these methods rely on specific noise assumptions or bandit structures, limiting their applicability to general settings. The recent work (Huang et al., 2023) develops a soft truncation method via the adaptive Huber regression to address these limitations. However, their method suffers undesired computational costs: it requires storing all historical data and performing a full pass over these data at each round. In this paper, we propose a one-pass algorithm based on the online mirror descent framework. Our method updates using only current data at each round, reducing the per-round computational cost from $\mathcal{O}(t\log T)$ to $\mathcal{O}(1)$ with respect to current round $t$ and the time horizon $T$, and achieves a near-optimal and variance-aware regret of order $\widetilde{\mathcal{O}}\big{(}dT^{\frac{1-\varepsilon}{2(1+\varepsilon)}}\sqrt{\sum_{t=1}^{T}\nu_{t}^{2}}+dT^{\frac{1-\varepsilon}{2(1+\varepsilon)}}\big{)}$ where $d$ is the dimension and $\nu_{t}^{1+\varepsilon}$ is the $(1+\varepsilon)$-th central moment of reward at round $t$.
Machine Learning, ICML
@@ 1 Introduction
Stochastic Linear Bandits (SLB) models sequential decision-making process with linear structured reward distributions, which has been extensively studied in the literature (Dani et al., 2008; Abbasi-Yadkori et al., 2011; Li et al., 2021b). Specifically, in SLB the observed reward at time $t$ is the inner product of the arm’s feature vector $X_{t}$ and an unknown parameter $\theta_{*}$, corrupted by sub-Gaussian noise $\eta_{t}$, namely,
@@
However, in many real-world scenarios, the noise in data may follow a heavy-tailed distribution, such as in financial markets (Hull, 2012) and online advertising (Jebarajakirthy et al., 2021). This motivates the study of heavy-tailed linear bandits (HvtLB) (Medina & Yang, 2016; Shao et al., 2018; Xue et al., 2020; Huang et al., 2023), where the noise $\eta_{t}$ in (1) satisfies that for some $\varepsilon\in(0,1],~{}\nu>0$,
@@
To address heavy-tailed noise, statistical methods in estimation and regression often rely on two key principles: truncation and median-of-means (MOM) (Lugosi & Mendelson, 2019). Truncation methods handle outliers by directly removing extreme data points. Soft version truncation, such as Catoni’s M-estimator (Catoni, 2012), reduces the impact of outliers by assigning them lower weights. This ensures robust estimates while maintaining potentially valuable information from extreme data points. Median-of-means (MOM) methods take a different approach by dividing the dataset into several groups, calculating the mean within each group, and then using the median of these group means as the final estimate. This method limits the influence of outliers to only a few groups, preventing impact on the overall dataset.
Table 1: Comparisons of our regret bounds and computational complexity to previous best-known results for heavy-tailed linear bandits. For the regret, the logarithmic dependence over $T$ is hidden by $\widetilde{\mathcal{O}}(\cdot)$-notation; $d$ is the dimension and $\nu_{t}$ is the moment bound. For the computational cost, we only keep the dependence on the time step $t$ and overall time step $T$.
@@
Huber HEAVY-OFUL (Huang et al., 2023) $\widetilde{\mathcal{O}}\left(dT^{\frac{1-\varepsilon}{2(1+\varepsilon)}}\sqrt{\sum_{t=1}^{T}\nu_{t}^{2}}+dT^{\frac{1-\varepsilon}{2(1+\varepsilon)}}\right)$ $\mathcal{O}(t\log T)$ instance-dependent bound
@@
For HvtLB, Shao et al. (2018) developed algorithms using truncation and MOM-based least squares approach for parameter estimation, and achieved an $\widetilde{\mathcal{O}}\big{(}dT^{\frac{1}{1+\varepsilon}}\big{)}$ regret bound which is proven to be optimal. However, these approaches have some notable issues: truncation methods require the assumption of bounded absolute moments for rewards, which cannot vanish in deterministic case, making them suboptimal in noiseless scenarios; MOM methods rely on repeated arm pulling and assumption of a fixed arm set, which makes them difficult to extend to more general settings. These limitations reveal that methods based on truncation and MOM heavily depend on specific assumptions or the bandit structure. For broad decision-making scenarios, such as online MDPs and adaptive control, where the state or context may vary dynamically, these assumptions no longer hold. A more detailed discussion of these challenges is provided in Section 6.
Recently, Sun et al. (2020) made remarkable progress in the study of heavy-tailed statistics by proposing adaptive Huber regression, which leverages the Huber loss (Huber, 1964) to achieve non-asymptotic guarantees. Li & Sun (2024) further applied this technique to SLB, achieving an optimal and variance-aware regret bound under bounded variance conditions ($\varepsilon=1$). Subsequently, Huang et al. (2023) extended the adaptive Huber regression to handle heavy-tailed scenarios ($\varepsilon\in(0,1]$), achieving an optimal and instance-dependent regret bound of $\widetilde{\mathcal{O}}\big{(}dT^{\frac{1-\varepsilon}{2(1+\varepsilon)}}\sqrt{\sum_{t=1}^{T}\nu_{t}^{2}}+dT^{\frac{1-\varepsilon}{2(1+\varepsilon)}}\big{)}$ for HvtLB. Notably, the adaptive Huber regression does not rely on additional noise assumptions or the repeated pulling of bandit settings. This generality allows the approach to be further adapted to other decision-making scenarios, such as reinforcement learning with function approximation.
@@
where $\Theta$ is the feasible set of the parameter, $\lambda$ is the regularization parameter, and $\ell_{s}(\theta)$ represents the Huber loss at round $s$. Unlike the OFUL algorithm (Abbasi-Yadkori et al., 2011) for SLB with sub-Gaussian noise, which uses least squares (LS) for estimation and can update recursively based on the closed-form solution, the Huber loss is partially quadratic and partially linear. Solving (3) requires storing all historical data and performing a full pass over all historical data at each round $t$ to update the parameter estimation. This results in a per-round storage cost of $\mathcal{O}(t)$ and computational cost of $\mathcal{O}(t\log T)$ with respect to the current round $t$ and the time horizon $T$, making it computationally infeasible for large-scale or real-time applications. Thus, we require a one-pass algorithm for HvtLB that updates the estimation at each round $t$ using only the current data, without storing historical data. Indeed, our goal is to design an algorithm for heavy-tailed linear bandits that is (basically) as efficient as OFUL in SLB with sub-Gaussian noise. Note that while one-pass algorithms based on the truncation and MOM approach have been proposed by Xue et al. (2023), but as noted earlier, these methods heavily rely on specific assumptions or the bandit structure. Given the attractive properties of Huber loss-based method (Huang et al., 2023), a natural question arises: Is it possible to design a one-pass Huber-loss-based algorithm that achieves optimal regret?
@@
Our Results.   In this work, we propose a Huber loss-based one-pass algorithm Hvt-UCB utilizing the Online Mirror Descent (OMD) framework, which achieves the optimal regret bound of $\widetilde{\mathcal{O}}(dT^{\frac{1}{1+\varepsilon}})$ while eliminating the need to store historical data. Additionally, our algorithm can further achieve the instance-dependent regret bound of $\widetilde{\mathcal{O}}\big{(}dT^{\frac{1-\varepsilon}{2(1+\varepsilon)}}\sqrt{\sum_{t=1}^{T}\nu_{t}^{2}}+dT^{\frac{1-\varepsilon}{2(1+\varepsilon)}}\big{)}$, where $\nu_{t}^{1+\varepsilon}$ is the $(1+\varepsilon)$-th central moment of reward at round $t$. Our approach preserves both optimal and instance-dependent regret guarantees achieved by Huang et al. (2023) while significantly reducing the per-round computational cost from $\mathcal{O}(t\log T)$ to $\mathcal{O}(1)$ with respect to the current round $t$ and the time horizon $T$. Moreover, our algorithm does not rely on additional assumptions, making it more broadly applicable. A comprehensive comparison of our results and previous works on heavy-tailed linear bandits is presented in Table 1.
@@
Key Challenges and Techniques.   Designing a one-pass algorithm for heavy-tailed linear bandits presents two major challenges. On the one hand, unlike the full-batch estimator (3(Huang et al., 2023), which leverages the entire historical data to control the estimation error, a one-pass estimator updates relies only on the current data point. Achieving comparable estimation error bound for one-pass estimator is much more challenging, especially in the presence of heavy-tailed noise. On the other hand, OMD is originally developed for regret minimization. Adapting it to parameter estimation requires linking the parameter gap (estimation error) with the loss value gap (regret). To address these challenges, our algorithm design and analysis incorporate two key components: (i) Handling heavy-tailed noise:  Zhang et al. (2016) have adopted a variant of Online Newton Step (ONS) to Logistic Bandits settings, but their method cannot be applied to HvtLB since their method relies on the global strong convexity of the logistic loss, whereas the Huber loss is partially linear. We generalize their approach to the OMD framework and introduce a recursive normalization factor that ensures normal data stays in the quadratic region of the Huber loss with high probability, maintaining robustness in estimation. This normalization factor also introduces a negative term in the estimation error, which is crucial for controlling stability. (ii) Supporting one-pass updates: The OMD framework introduces a stability term in estimation error, which represents the impact of online updates. The adaptive normalization factor adds a multiplicative bias to this term, resulting in a positive term of order $\widetilde{\mathcal{O}}(\sqrt{T})$ that weakens the bound. This requires a careful design of the normalization factor and learning rate of OMD to cancel it out using the negative term that arises from our OMD-based analysis, which is similar to a crucial technique in the regret analysis of recent gradient-variation online learning methods (Zhao et al., 2020b, 2024).
@@ 2 Related Work
Heavy-Tailed Bandits.   The multi-armed bandits (MAB) problem with heavy-tailed rewards, characterized by finite $(1+\varepsilon)$-moment, was first studied by Bubeck et al. (2013). They introduced two widely used robust estimation and regression methods from statistics into the heavy-tailed bandits setting: truncation and median-of-means (MOM). Medina & Yang (2016) extended these methods to heavy-tailed linear bandits (HvtLB) and proposed two algorithms: the truncation-based CRT algorithm and the MOM-based MoM algorithm, achieving $\widetilde{\mathcal{O}}\big{(}dT^{\frac{2+\varepsilon}{2(1+\varepsilon)}}\big{)}$ and $\widetilde{\mathcal{O}}\big{(}dT^{\frac{1+2\varepsilon}{1+3\varepsilon}}\big{)}$ regret bounds, respectively. Later, Shao et al. (2018) established an $\Omega\big{(}dT^{\frac{1}{1+\varepsilon}}\big{)}$ lower bound for HvtLB and introduced the truncation-based TOFU algorithm and the MOM-based MENU algorithm, both achieving an $\widetilde{\mathcal{O}}\big{(}dT^{\frac{1}{1+\varepsilon}}\big{)}$ regret bound. For HvtLB with finite arm set, Xue et al. (2020) established an $\Omega\big{(}d^{\frac{\varepsilon}{1+\varepsilon}}T^{\frac{1}{1+\varepsilon}}\big{)}$ lower bound, and proposed the truncation-based BTC algorithm and MOM-based BMM algorithm, both of which achieve an $\widetilde{\mathcal{O}}\big{(}\sqrt{d}T^{\frac{1}{1+\varepsilon}}\big{)}$ regret bound, reducing the dependence on the dimension $d$ under the finite arm set setting. To the best of our knowledge, the two works (Kang & Kim, 2023; Li & Sun, 2024) first introduce Huber loss to linear bandits. Specifically, Kang & Kim (2023) studied the heavy-tailed linear contextual bandits (HvtLCB) with fixed and finite arm set, where the context $X_{t}$ at each round is independently and identically distributed from a fixed distribution. They proposed the Huber-Bandit algorithm based on Huber regression, achieving an $\widetilde{\mathcal{O}}\big{(}\sqrt{d}T^{\frac{1}{1+\varepsilon}}\big{)}$ regret bound, On the other hand, Li & Sun (2024) were the first to apply adaptive Huber regression technique (Sun et al., 2020) to SLB, focusing on noise with finite variance (namely, HvtLB with $\varepsilon=1$). They proposed the AdaOFUL algorithm, achieving an optimal variance-aware regret of $\widetilde{\mathcal{O}}(d\sqrt{\sum_{t=1}^{T}\nu_{t}^{2}})$, where $\nu_{t}$ is the upper bound of the variance. This result is subsequently improved by Huang et al. (2023), which is also the most related work to ours. Huang et al. (2023) considered general heavy-tailed scenarios with $\varepsilon\in(0,1]$ and proposed HEAVY-OFUL with an instance-dependent regret of $\widetilde{\mathcal{O}}\big{(}dT^{\frac{1-\varepsilon}{2(1+\varepsilon)}}\sqrt{\sum_{t=1}^{T}\nu_{t}^{2}}+dT^{\frac{1-\varepsilon}{2(1+\varepsilon)}}\big{)}$ where $\nu_{t}^{1+\varepsilon}$ is the $(1+\varepsilon)$-th central moment of reward at round $t$. Subsequent works have applied Huber regression to various bandit settings, such as matrix-valued contextual bandits with low-rank structure (Kang et al., 2024). We emphasize that all the above Huber regression-based algorithms for heavy-tailed bandits are not updated in a one-pass manner. Beyond Huber-based approaches, other techniques have been explored for handling heavy-tailed noise and outliers. These include general robust loss functions like Tilted Empirical Risk Minimization (TERM) (Li et al., 2021a), and reweighting-based methods proposed in the offline contextual bandit (Sakhi et al., 2024; Behnamnia et al., 2024). A potential future direction could be to develop one-pass algorithms based on these techniques.
@@
One-Pass Bandit Learning.   In SLB, Abbasi-Yadkori et al. (2011) proposed the OFUL algorithm, achieving an $\mathcal{O}(d\sqrt{T})$ regret bound. It uses Least Squares (LS) for parameter estimation, naturally supporting one-pass updates with a closed-form solution. As an important extension of SLB, Generalized Linear Bandits (GLB) is first introduced by Filippi et al. (2010). Their proposed GLM-UCB algorithm achieves an $\mathcal{O}(\kappa d\sqrt{T})$ regret bound, where $\kappa$ measures the non-linearity of the generalized linear model. This algorithm relies on Maximum Likelihood Estimation (MLE) which does not have a closed-form solution like LS in SLB and requires storing all historical data, resulting in a per-round computational cost of $\mathcal{O}(t)$. To this end, Zhang et al. (2016) proposed OL2M, the first one-pass algorithm for Logistic Bandits (a specific class of GLB), achieving an $\mathcal{O}(\kappa d\sqrt{T})$ regret bound with per-round computational cost of $\mathcal{O}(1)$. Later, Jun et al. (2017) explored one-pass algorithms for GLB by introducing the online-to-confidence-set conversion. Using this conversion, they developed GLOC, a one-pass algorithm that achieves the same theoretical guarantees and per-round cost as OL2M but is applicable to all GLB. Xue et al. (2023) considered heavy-tailed GLB and proposed two one-pass algorithms: truncation-based CRTM and mean-of-medians-based CRMM, both achieving an $\widetilde{\mathcal{O}}\big{(}\kappa dT^{\frac{1}{1+\varepsilon}}\big{)}$ regret bound. Subsequent works have tackled more challenging topics, such as reducing the dependence on $\kappa$ with one-pass algorithms or designing such algorithms for more complex models like Multinomial Logistic Bandits and Multinomial Logit MDPs (Zhang & Sugiyama, 2023; Li et al., 2024; Lee & Oh, 2024; Li et al., 2025).